# Circuit Component Reuse Across Tasks in Transformer Language Models

### v1

@@ 1 Introduction
To better understand this, we study whether LMs reuse model components across different tasks to accomplish the same general behaviors, and whether these compose together in circuits in predictable ways. We study two tasks that have no obvious linguistic overlap but that we hypothesize may require similar processes to solve. One is the Indirect Object Identification (IOI) task, for which Wang et al. (2022) discover a circuit in GPT2-Small (Radford et al., ). The other is the Colored Objects task (Figure 1), which is a variation on an in-context learning task from BIG-Bench (Ippolito & Callison-Burch, 2023). To solve the Colored Objects task, a model must copy a token from a list of possible options in context. Based on the interpretation of the IOI circuit, which performs a behavior along those lines, a simple way to test the idea of neural reuse would be to see if these two tasks use the same circuit. We perform path patching (Wang et al., 2022; Goldowsky-Dill et al., 2023) on both of these tasks on GPT2-Medium (Sections 3 and 4) and compare their circuits. Our causal interventions show that the Colored Objects circuit uses largely the same process, and around 78% of the ‘in-circuit’ attention heads are shared between the two tasks [^1] Quantifying overlap is a difficult problem that we do not have an exact solution for. We discuss this in Appendix H.. Such a high degree of overlap supports the idea of circuit general reuse. We use this insight to design ‘ideal’ interventions on the model that both improve the Colored Objects performance from 49.6% to 93.7% and provide evidence that at least part of the original IOI circuit appears to be part of a more generic circuit for controlling copying a token from possible choices in context.
@@ 4 The Colored Objects Circuit
The above process is remarkably similar algorithmically to that used for IOI: the detection of duplication informs the mover heads either where or where not to look. Figure 3 visualizes this overlap and the relative importance of each head, which shows a very large overlap between the heads performing this function; thresholding at the 2% most important heads for each circuit, we find that 25/32, or 78% of the circuit is shared. This is a difference of swapping out three inhibition heads and a negative mover head for three content gatherer heads. It is worth noting that such overlap is not trivial–i.e., it is not as though any two circuits will contain the same components or even overlapping algorithmic steps. See, for example, the circuit described for the greater-than task in Hanna et al. (2023), which looks fundamentally different333compared to IOI on GPT2-Small there is essentially no reported overlap in specific behaviors and heads, but it is difficult to say without comparing the two circuits directly. In the following sections, we will describe the evidence for the roles we prescribe to components in circuit, before returning to evaluating the differences to the IOI circuit. For consistency with the path patching process, we decribe components working backward from the end of the algorithm to the beginning.
@@ 4.2 Content Gatherer Heads
To confirm that they are playing this role, we run an intervention as follows: at the [end] position, we block attention to the words in the question part of the test prompt about which we hypothesize the CGs are passing information. In particular, if these heads account for all of the signal that tells the model which object is being asked about, then the mover heads should randomly copy between the three color words in the test example. By doing so, accuracy is reduced to 35%, i.e., the model is selecting the correct color from among the three test example colors roughly at random chance. Further analysis on these heads is done in Appendix C.2.
@@ B.2 The Negative Mover Head
Figure 8: Color bar shows percent logit difference as a result of the patching.
@@ Appendix G Details on Colored Objects Dataset Construction
Appendix H Quantifying Circuit Overlap
@@
Putting an exact number on how similar two circuits are is a very complex task that we do not attempt to fully answer in this paper. At the moment, the criteria for deciding whether any component is considered in the circuit or not is not very well defined, although there are best practices defined by previous circuit analyses (Wang et al., 2022; Lieberum et al., 2023; Goldowsky-Dill et al., 2023; Hanna et al., 2023) and there is some work exploring automating this process (Conmy et al., 2023). Thus, we do our best to place an accurate number on the overlap as the proportion of shared nodes in the circuit. We study three processing ‘steps’ in each circuit which also makes the job of assigning similarity easier, but it could be possible that these numbers differ between two tasks for the purpose of other analyses. Our results motivate studying the problem of quantifying circuit overlap in future work.
Appendix I How was the Colored Objects Task Selected?
@@
Appendix J Do Other Models also Use Similar Circuits for These Tasks?
We perform a preliminary path analysis of larger versions of GPT2 (Large and XL) by path patching from attention heads to the logits for the IOI and Colored Objects tasks (Figure 23). We find that there isn’t an obvious overlap between the heads that most directly impact the logits in either model, however it does appear that mover heads are playing a role in both cases. This also shows that path patching would not just assign the same importance for certain nodes across these two tasks.
Figure 23: Path patching from attention heads to the logits of GPT2-Large and GPT2-XL show no clear overlap like the one we saw in GPT2-Medium with Colored Objects and IOI. However, due to the number of mover heads in the larger models, it’s possible this would be less likely overall, and the underlying circuits may still be similar functionally. Colored bars show percent logit difference.

### v2

@@ 1 Introduction
To better understand this, we study whether LMs reuse model components across different tasks to accomplish the same general behaviors, and whether these compose together in circuits in predictable ways. We study two tasks that have no obvious linguistic overlap but that we hypothesize may require similar processes to solve. One is the Indirect Object Identification (IOI) task, for which Wang et al. (2022) discover a circuit in GPT2-Small (Radford et al., ). The other is the Colored Objects task (Figure 1), which is a variation on an in-context learning task from BIG-Bench (Ippolito & Callison-Burch, 2023). To solve the Colored Objects task, a model must copy a token from a list of possible options in context. Based on the interpretation of the IOI circuit, which performs a behavior along those lines, a simple way to test the idea of neural reuse would be to see if these two tasks use the same circuit. We perform path patching (Wang et al., 2022; Goldowsky-Dill et al., 2023) on both of these tasks on GPT2-Medium (Sections 3 and 4) and compare their circuits. Our causal interventions show that the Colored Objects circuit uses largely the same process, and around 78% of the ‘in-circuit’ attention heads are shared between the two tasks [^1] Quantifying overlap is a difficult problem that we do not have an exact solution for. We discuss this in Appendix H.3.. Such a high degree of overlap supports the idea of circuit general reuse. We use this insight to design ‘ideal’ interventions on the model that both improve the Colored Objects performance from 49.6% to 93.7% and provide evidence that at least part of the original IOI circuit appears to be part of a more generic circuit for controlling copying a token from possible choices in context.
@@ 4 The Colored Objects Circuit
The above process is remarkably similar algorithmically to that used for IOI: the detection of duplication informs the mover heads either where or where not to look. Figure 3 visualizes this overlap and the relative importance of each head, which shows a very large overlap between the heads performing this function; thresholding at the 2% most important heads for each circuit, we find that 25/32, or 78% of the circuit is shared. This is a difference of swapping out three inhibition heads and a negative mover head for three content gatherer heads. It is worth noting that such overlap is not trivial–i.e., it is not as though any two circuits will contain the same components or even overlapping algorithmic steps. See, for example, the circuit described for the greater than task in Hanna et al. (2023), which looks fundamentally different333compared to IOI on GPT2-Small there is essentially no overlap in specific behaviors and heads. In the following sections, we will describe the evidence for the roles we prescribe to components in circuit, before returning to evaluating the differences to the IOI circuit. For consistency with the path patching process, we decribe components working backward from the end of the algorithm to the beginning.
@@ 4.2 Content Gatherer Heads
To confirm that they are playing this role, we run an intervention as follows: at the [end] position, we block attention to the words in the question part of the test prompt about which we hypothesize the CGs are passing information. In particular, if these heads account for all of the signal that tells the model which object is being asked about, then the mover heads should randomly copy between the three color words in the test example. By doing so, accuracy is reduced to 35%, i.e., the model is selecting the correct color from among the three test example colors roughly at random chance. For three seeds, we randomly sample three heads from these layers (without replacement) that are not predicted as playing a role in content-gathering and repeat this intervention the average accuracy for random heads is 49.9%±$\pm 1.0$. Further analysis on these heads is done in Appendix C.2.
@@ 7.3 Why can’t GPT2 activate the Inhibition Heads on its own?
8 Acknowledgments
We would like to thank Aaron Traylor, Catherine Chen, Michael Lepori, Etha Hua, and members of the Google Deepmind mechanistic interpretability team for feedback on this work.
@@ B.2 The Negative Mover Head
Figure 8: Path patching to the queries of the negative mover head 19.1. Color bar shows percent logit difference as a result of the patching. This head relies on a different set of heads than the positive mover heads.
@@ Appendix G Details on Colored Objects Dataset Construction
Appendix H Circuit Overlap in Additional Tasks
Comparing the IOI and Colored Objects tasks present an interesting case study that lets us highlight two different tasks with very similar circuits. In this section, we analyze two additional tasks that vary in their similarity to these two tasks to provide additional reference points for quantifying similarity. The two new tasks are the Greater Than dataset (Hanna et al., 2023) and a simple factual recall task in which the model predicts the capital city of a country. While both tasks are fundamentally different from the IOI and Colored Objects tasks, prior work suggests specific hypotheses about the amount overlap we can expect in these circuits.
Greater Than Dataset
In this task, originally introduced in Hanna et al. (2023), the model must predict a viable end year for some fictional event. For example: “The war lasted from the year 1732 to the year 17”, where any number between 33-99 is possible, but 00-32 is not. The authors analyze the circuit GPT2-Small uses to solve this task, which makes significant use of the MLP layers. Therefore, we expect the circuit GPT2-Medium uses to have very little overlap with the one used for IOI or Colored Objects. We use the code provided by Hanna et al. (2023) to generate a dataset of 200 examples to analyze.
World Capitals Dataset
Given a sentence like “The capital of France is” the model must access information that is not provided in context to predict the right city. This makes the task different from either the IOI and Colored Objects tasks. However, Geva et al. (2023) provide an interpretation for how such factual recall happens in LMs (not through circuit analysis). In their analysis, early MLPs enrich the subject token (the “France” token), and later attention heads extract the capital city attribute by attending from “is” to “France”. In attribute extraction, some attention heads attend to subject tokens and copy information out of them. If this is the process that is used for this task, our earlier results might predict that the mover heads are reused to perform this operation. We perform path patching on a small dataset 200 pairs of different countries to reverse engineer the capital city prediction process. We filter out countries and capitals that are not tokenized as a single token by GPT2-Medium, giving us 47 unique countries to generate the data from.
H.1 Greater Than Circuit
Figure 23: Heatmaps from path patching results from reverse engineering the Greater Than task from (Hanna et al., 2023) on GPT2-Medium. The rightmost column represent the MLPs in each layer.
The results of our path patching experiments on this dataset are shown in Figure 23. We path patch using the probability difference metric for all years (see Hanna et al. (2023) for details), which is analogous to the logit difference metric. We first path patch to the logits from the end position and find the MLPs are most important. From there we path patch to last token position to these MLPs, and then backwards through the last token position from there. The circuit can be summarized as follows:
Attention heads 6.1, 5.8, and 6.15, write into 7.2, 9.9, 11.1, and 11.5, 13.12, 14.14, MLPs 14 and 15. In turn these latter components write into MLPs 17-23. To summarize, the earlier induction heads (7.2, 9.9, 11.1, etc. see Figure 17) attend from the last token (e.g., “17”) to the last two digits in the start year (e.g., the “32” in “1732”). 14.14 is a significant mover head in the Colored Objects task and has the same pattern. Similarly to Hanna et al. (2023), the MLPs perform the majority of the numerical comparison needed to predict a reasonable year. As expected, this leads to a very low overlap with the two tasks we study in the main paper, in which the MLPs play almost no part.
Summary of Overlap:
The Greater Than task represents a task with very low overlap, but there are still components shared with the IOI and Colored Objects circuits. Induction heads 6.15, 7.2, and 9.9 also contribute significantly in IOI and Colored Objects, but this is due to simply matching an induction pattern (Olsson et al., 2022). Mover head 14.14 plays a similar role in Colored Objects. If we just consider the 18 components in the Greater Than circuit, the Colored Objects task has an overlap of 4/18=22.2% and IOI has an overlap of 3/18=16.7%
H.2 World Capitals Circuit:
Figure 24: World Capitals path patching results; the rightmost column in the heatmaps represent the MLPs of the layers. Given inputs such as ”The capital of France is”, the majority of behavior was captured in two steps. First, mover heads attend from the final token position to the country and extract the capital city. The value vectors of the country position for these heads depend on the MLPs in the earlier layers (in particular, MLP 0).
The path patching results for the World Capitals task are shown in Figure 24. Given prompts of the form ”The capital of France is”, we first path patch to the logits at the last token position (“is”), and find that the majority of components that write to the logits are mover heads. We path patch to the queries, keys, and values of these heads and find that the values at the country position have the greatest effect on the output of the mover heads, which are predominantly dependent on the MLPs in the early layers. The circuit is summarized below:
MLP 0, 1, 6, and 9-14 build up the representation at the country position of the input. Attention heads 14.7, 15.0, and the mover heads 16.2, 19.15, 20.6, 20.12, and 21.12 attend to the country token and promote the capital city in the logits.
Summary of Overlap:
Despite the differences between this task and the Colored Objects and IOI tasks, there is significant overlap, particularly with the use of mover heads to copy the final answer. The mover heads that overlap with Colored Objects and IOI are 16.2, 19.15, 20.6, and 20.12. Therefore, 4/7 of the mover heads are shared. Although less prominent, 17.4 and 15.14 also contribute to the world capitals logits which appeared previously as mover heads (Figure 4.1). This aligns with our hypothesis that mover heads would be used for moving information about the capital out of the country. The rest of the circuit relies heavily on the MLPs, so only 4/16=25% of the full circuit components overlap with both the Colored Objects and IOI tasks, demonstrating how different the earlier processing is.
H.3 Challenges in Quantifying Circuit Overlap
@@
Putting an exact number on how similar two circuits are is a very complex task that we do not attempt to fully answer in this paper. At the moment, the criteria for deciding whether any component is considered in the circuit or not is not very well defined, although there are best practices defined by previous circuit analyses (Wang et al., 2022; Lieberum et al., 2023; Goldowsky-Dill et al., 2023; Hanna et al., 2023) and there is some work exploring automating this process (Conmy et al., 2023). Thus, we do our best to place an accurate number on the overlap as the proportion of shared nodes in the circuit. In the main paper, we study three processing ‘steps’ in each circuit which also makes the job of assigning similarity easier, but it could be possible that these numbers differ between two tasks for the purpose of other analyses. Our results motivate studying the problem of quantifying circuit overlap in future work.
Appendix I Do Other Models also Use Similar Circuits for These Tasks?
We perform a preliminary path analysis of larger versions of GPT2 (Large and XL) by path patching from attention heads to the logits for the IOI and Colored Objects tasks (Figure 25). We find that the overlap between the heads used in either task is not as obvious in either model, however it does appear that mover heads are playing a role in both cases. For both models, we compare the top 10 heads in the Colored Objects and IOI tasks. In GPT2-Large, five of these heads overlap, but none overlap in GPT2-XL. These heads are as follows:
GPT2-Large (Overlapping heads bolded):
IOI: 20.14, 30.13, 27.17, 29.0, 23.13, 26.14, 22.0, 32.18, 30.10, 21.8
Colored Objects: 22.0, 23.13, 23.7, 25.13, 21.8, 21.14, 26.14, 24.13, 31.9, 27.17
GPT2-XL:
IOI: 41.9, 36.17, 42.7, 28.22, 40.4, 39.9, 39.4, 35.17, 37.15, 38.12
Colored Objects: 24.16, 26.20, 28.15, 21.13, 40.18, 29.5, 30.21, 33.18, 31.3, 42.9
It appears that the overlap in the top most important heads decreases as the scale of the model goes up from GPT2-Medium to GPT2-XL, however, we find that the functions of the heads stays roughly the same, meaning that many of these heads are mover heads. We measure this by characterizing the function of each head identified in the IOI circuit on the Colored Objects task. This tells us whether the heads are performing similar roles, even if they don’t light up as the most important according to path patching results. In GPT2-Large, the 5 shared heads are all mover heads (21.8, 22.0, 23.13, 26.14, 27.17). In GPT2-XL, even though the exact heads being used are different, 4 of the heads are mover heads (35.17, 36.17, 38.12, 39.4) and 1 is a negative mover head (39.9) on both tasks. As models get larger, there are more ‘paths’ through the network, which makes it less likely that exactly the same heads are used for different tasks, but it appears there is still overlap in the function of the important heads. This is important to keep in mind, as it complicates the picture for analyzing larger models, but the upshot appears that we can still expect some amount of functional overlap between similar tasks.
Figure 25: Path patching from attention heads to the logits of GPT2-Large and GPT2-XL. There is some overlap in GPT2-Large, and almost none in the most important heads for GPT2-XL, suggesting that as the number of components increases, overlap is less likely. However, we find that most of the important heads are mover heads in both cases, suggesting the underlying circuits may still be similar functionally even if they take different paths through the models. Colored bars show percent logit difference.
Appendix J How was the Colored Objects Task Selected?